6) You create 40 mL of a 3.0 M buffer with a pH of 5. You add 7 mL of a 1.25M H2SO4 solution to the buffer you made, what is the new pH?

7) You create 50 mL of a 1.0M buffer with a pH of 2. You add 2 mL of a .45M HCl solution to the buffer you made, what is the new pH?

8) You create 30 mL of a 1.0M buffer with a pH of 7.5. You add 1.3 mL of a .66M HCl solution to the buffer you made, what is the new pH?

9) You create 100 mL of a 0.5M buffer with a pH of 6.8. You add 1.0mL of a 1.0M NaOH solution to the buffer you made, what is the new pH?

10) You create 100 mL of a 0.5M buffer with a pH of 11.75. You add 10.0 mL of a 1.0M HCl solution to the buffer you made, what is the new buffer [Base]?
1
Expert's answer
2020-06-10T10:51:31-0400

6) C(acid) C(base)

Initial(I) (40 ml*3.0M /1000) 0.12 0.12

Change(C)(7ml*1.25 M/1000) +0.00875 -0.00875

Final(F) 0.12875 0.11125

pH=5-lg[Acid]/[base]= 5 - lg(0.12875/0.11125)=5-0.0634=4.9366

7) C(acid) C(base)

I 0.05 0.05

C +0.0009 -0.0009

F 0.0509 0.0491

pH=2-0.0156=1.984


8) C(acid) C(base)

I 0.03 0.03

C +0.000858 -0.000858

F 0.030858 0.029142

pH=7.5 - 0.0248=7.4752


9) C(Acid) C(Base)

I 0.05 0.05

C -0.001 +0.001

F 0.049 0.051

pH= 6.8 - (-0.0174)=6.8174


10) C(acid) C(base)

I 0.05 0.05

C +0.01 -0.01

F 0.06 0.04

pH=11.75-0.176=11.574